How many different groups of 3 can you make you make with 22 students?
Answer by jim_thompson5910(35256)

n C r = (n!)/(r!(n-r)!)

22 C 3 = (22!)/(3!*(22-3)!)

22 C 3 = (22!)/(3!*19!)

22 C 3 = (22*21*20*19!)/(3!*19!)

22 C 3 = (22*21*20)/(3!)

22 C 3 = (22*21*20)/(3*2*1)

22 C 3 = (9240)/(6)

22 C 3 = 1540

So there are 1,540 different groups (assuming order doesn't matter)